Consider a Poisson process with rate between the 5rd and 7th arrivals? = 3. What is the distribution of the time

Name the distribution and any parameters. Show mathematically, the distribution between the 5th and 7th arrival.
Answer #1

Note that poisson probability distribution and exponential distribution follows memoryless property. Therefore the amount of waiting time between 5th and 7th arrival is same as the waiting time for 2 arrivals. Let the number of arrivals be given by random variable X and waiting time by random variable Y.

Then the distribution of waiting time is given as:

P(Y > y) = Probability that waiting time is more than y

which should be equal to probability that there is no or 1 arrival in time y

P(Y > y) = P(X = 0) + P(X = 1)

Now for time y, the distribution for X is given as:

X Poisson(Ay)

X Poisson (3y)

Therefore the probability now is computed as:

P(YyP(X 0)P(X = 1)

P(Yy)e 3ye 3y

Therefore, we have the cumulative distribution of Y given as:

F(y)P(Y y) 1- P(Y> y) 1 -e-3ye3y

Therefore the distribution now can be obtained by differentiating the cumulative distribution with respect to y as:

9yey 9ye

This is the required distribution of the time here.